Find the conjugate of the complex number (1+i)^2n+1 / (1-i)^2n-1 , n is natural no.


 


 


 

Grade:10

1 Answers

Md Rizwan
11 years ago

(1+i)/(1-i)   = (1+i)2 /[(1+i)*(1-i)]   = 2i /2 =i
(1+i)2n+1 /(1-i)2n-1   = [(1+i)/(1-i)]2n *[(1+i)*(1-i)] = i2n * 2 = 2[(i)2]n= 2*(-1)n
so, it is purely real, its conjugate is the number itself.
hope you understood